OpenStudy (anonymous):

Find the general solution of tan2x = -cot(x+pie/3)?

OpenStudy (anonymous):

u kno tan2x = 2tanx/1-tan^2x nd cotx = 1/tanx

OpenStudy (anonymous):

form a quadratic nd solve........ for tanx........ then u l get the g.s

OpenStudy (anonymous):

\[\tan(2x) = -\cot(90 + x -30) \implies -(-\tan(x-30)) = \tan(x - \frac{\pi}{6})\] \[\implies \tan(2x) = \tan(x - \frac{\pi}{3})\]

OpenStudy (anonymous):

General solution is given by: \[\tan ( \theta) = \tan(x)\] then: \[\theta = n \pi + x\]

OpenStudy (anonymous):

so: \[2x = n \pi + ( x - \frac{\pi}{3} ) \implies x = n \pi + \frac{\pi}{3}\]

OpenStudy (anonymous):

hey @waterineyes but my books says it as x = n pie + 5pie/6

OpenStudy (anonymous):

So put n = 0 1 2 3 4 etc etc and find the corresponding value of x..

OpenStudy (anonymous):

you can subtract it from \(2 \pi\)..

OpenStudy (anonymous):

wat???

OpenStudy (anonymous):

Wait something got wrong.. Let me check..

OpenStudy (anonymous):

@w\ r uthere

OpenStudy (anonymous):

Wait I am checking it..

OpenStudy (anonymous):

I got it..

OpenStudy (anonymous):